AFM2TFM(1) AFM2TFM(1) NAME afm2tfm−convertAdobefontmetricstoTeXfontmetrics DESCRIPTION PostScriptfontsare(orshouldbe)accompaniedbyfontmetricfilessuchasTimes-Roman.afm, whichdescribesthecharacteristicsofthefontcalledTimes-Roman.TousesuchfontswithTeX, weneedTFMfilesthatcontainsimilarinformation.Thisprogramdoesthatconversion.Formore information,printoutthedvipsmanual. FILES psfonts.mapintheconfigdirectoryusedbydvips(updatethis). ../afm/Makefileinthedvipssources(lookatthisforexamples). SEEALSO dvips(1),tex(1),vptovf(1) BUGS AnoptiontoaddletterspacingtothevirtualfontwasrejectedbyKnuthbecauseitwouldmake thefiandflligaturesbreakthenormalrhythmicpattern. AUTHORS TomasRokicki<rokicki@cs.stanford.edu>andDonaldE.Knuth 9August1990 1